# Leveraging Label Information for Stealthy Data Stealing in Vertical Federated Learning

### v1

@@ 6.2. Baselines and defenses
We also implement SOTA defensive methods, SplitGuard (SG) and Gradient Scrutinizer (GS) on malicious attacks, to assess reconstruction performance under these defenses.
@@
Gradient Scrutinizer (GS): Evaluates the detection score of received gradients in each round to identify potential attacks. Due to instability in detection scores leading to frequent misjudgments, we adopt an average score strategy to improve reliability and set the threshold to 0.8 in our experiments. Similarly, when the average detection score falls below the detection threshold, an attack is detected.
@@
In our experimental results, we append the GS or SG to the names of baseline methods to indicate their evaluation under defense mechanisms, SG or GS. We add (GS/SG) to our methods indicating the evaluation of the attack with and without defenses, GS and SG.
6.3. Data reconstruction results
@@
Overview: As shown in Table 1, for all datasets, our proposed DMAVFL and DMAVFL_sync consistently outperform all other approaches in achieving the lowest reconstruction errors and embedding cosine distance. This underscores the effectiveness of our techniques and the advantage of embedding distribution transfer using DAC. For the RT_IOT2022 dataset, DMAVFL_sync achieves a reconstruction error of 1.2324 and cosine distance of 0.1025, significantly outperforming FSHA, which has a reconstruction error of 1.5150 and cosine distance of 0.4929.
@@
Our methods demonstrate robustness against defensive mechanisms, maintaining consistent performance and circumventing both tested defense methods for malicious attacks. In sharp contrast, the performance of AGN and FSHA become worse when faced with the defenses, and even completely failed under some circumstances. Specifically, AGN’s reconstruction error increases dramatically from 0.0929 to 0.5067, and FSHA’s from 0.0153 to 0.5029 for the MNIST dataset. This demonstrates the stealthiness and adaptability of DMAVFL and DMAVFL_sync compared to other methods.
@@
An important observation from our experiments is that a lower reconstruction error does not necessarily correspond to a smaller embedding distance. For example, for the MNIST dataset, while DMAVFL_sync achieves a lower reconstruction error than DMAVFL, it exhibits a slightly higher MSE and cosine embedding distance. This can be attributed to the synchronous training of all model components in DMAVFL_sync, which promotes dynamic learning and results in model diversity. Despite the higher average embedding distance, the decoder in DMAVFL_sync is more effective at reconstructing features, effectively compensating for the increased distance. Additionally, a smaller MSE embedding distance does not guarantee a more accurate reconstruction or a more similar embedding structure, as evidenced by GIA in the Credit dataset and FSHA in the MNIST dataset, which has the smallest MSE embedding distance but significantly higher cosine embedding distance and reconstruction error compared to other methods.
@@
We do not record AGN-GS and FSHA-GS in the Credit dataset. It is because GS doesn’t work in the Credit dataset, which can also be observed in Figure 6(h), caused by the limited data features (23 features) and imbalanced label distribution (77% label 0 and 23% label 1).
@@
Visualization of the reconstruction: Table 2 visualizes the reconstruction performance of the proposed DMAVFL, alongside the malicious attack baselines, AGN and FSHA, for the CIFAR-10 dataset. The results clearly demonstrate that DMAVFL produces more accurate and distinct reconstructions compared to the baselines. While AGN and FSHA yield coarser results, DMAVFL maintains high-quality reconstruction even under defensive conditions. In contrast, both FSHA and AGN struggle to reconstruct informative pixels effectively when faced with defense mechanisms.
@@
Table 2. Visualization of reconstructed features on the CIFAR-10 dataset with and without the defense. Except for the original image, the right side of each image is the reconstructed image. The last column is the average reconstruction error.
@@
SG and GS defense results: Figure 6 illustrates the detection scores for DMAVFL and DMAVFL_sync, which consistently exceed the detection thresholds, aligning indistinguishably with honest training scenarios. Although in the CIFAR-10 dataset, DMAVFL’s detection score is lower than the threshold at the beginning, its average score remains above the threshold, ensuring stealthiness against target clients. In contrast, malicious attack baselines, AGN and FSHA, consistently maintain scores below the detection thresholds of SG and GS in the MNIST and CIFAR-10 datasets, rendering them detectable.
@@
(c) SG scores in RT_IOT2022.
(d) SG scores in Credit.
(e) GS scores in MNIST.
(f) GS scores in CIFAR-10.
(g) GS scores in RT_IOT2022.
(h) GS scores in Credit.
Figure 6. Detection scores of SG and GS on malicious attacks and the honest training across four datasets.
7. Ablation Study
This section explores the robustness of DMAVFL and DMAVFL_sync (abbreviated as sync) by evaluating their performance under various parameter settings. These settings include changes in the adversary’s feature size, use of heterogeneous encoder models, variations in auxiliary dataset size, and differences in the number of target clients. Except for the variables specifically altered for this study, the model setups and implementation environment remain consistent with those described in the main experiments.
7.1. Effect of adversary’s feature size
To explore how the size of the feature set held by the adversary affects the reconstruction performance of DMAVFL and sync, we conduct experiments targeting the reconstruction of 30% of the features while varying the proportion of features held by the adversary from 0% to 70%. At 0% feature possession, the scenario reduces to split learning where no features or bottom model are held by the adversary. These experiments are performed on the Credit and RT_IOT2022 (abbreviated as IOT) datasets, with the results summarized in Table 3.
Table 3. Reconstruction error under various adversary’s feature portions.
Feature portion 0% 10% 20% 50% 70%
Credit DMAVFL 0.8315 0.7839 0.7489 0.6987 0.5189
sync 0.8718 0.7537 0.7154 0.6940 0.5732
IOT DMAVFL 1.7423 1.7014 1.6939 1.6382 1.5456
sync 1.7848 1.6721 1.6103 1.6056 1.2811
Consistent with our expectations, increasing the proportion of features held by the adversary leads to improved reconstruction accuracy. As illustrated in Table 3, the reconstruction error in the Credit dataset using DMAVFL decreases from 0.8315 to 0.5189 as the feature possession increases from 0% to 70%. This trend is observed for both datasets and methods.
7.2. Effect of encoder’s complexity
@@
Table 4. Reconstruction error and average embeddings distance using different encoder’s model.
Encoder Metric DMAVFL sync
@@
Table 4 clearly illustrates that the adversary achieves the lowest reconstruction error and the smallest embedding cosine distance when using an encoder the same as the target model, as observed for both DMAVFL and sync. This optimal alignment underscores the importance of model congruence in enhancing the effectiveness of embedding distribution transfer and data reconstruction. Conversely, when the encoder structure deviates from that of the target model, specifically when switching from a CNN to an MLP, the reconstruction performance significantly deteriorates, with reconstruction error increasing by an order of magnitude (from 0.0295 to 0.2545). This dramatic drop highlights the challenges of using less compatible models for data reconstruction attack. Moreover, the results indicate that a complex encoder, featuring a higher number of residual blocks, outperforms a simpler encoder configuration. The complex encoder not only enhances the decoder’s ability to reconstruct the data accurately but also learns a more informative embedding. This suggests that increasing the encoder’s complexity, when the target model is unknown, can lead to better performance in both embedding distribution transfer and data reconstruction.
7.3. Effect of auxiliary dataset
@@
In the main experiments, we set the size of the adversary’s auxiliary dataset, $|D_{aux}|$, relative to the training set size, $|D_{train}|$, at a ratio of $1:10$. To explore the impact of varying auxiliary dataset sizes, we adjust the ratio of $|D_{aux}|:|D_{train}|$ from $0:10$ to $1:1$ in the MNIST dataset. Specifically, at a ratio of 0:10, the auxiliary dataset comprises the FashionMNIST (Xiao et al., 2017) training dataset instead of a subset of MNIST, introducing a distinct distribution from the training set. This setup aims to test the adaptability of the reconstruction process to changes in the auxiliary dataset’s distribution.
Table 5. Reconstruction error for different $|D_{aux}|$ relative to $|D_{train}|$.
$|D_{aux}|:|D_{train}|$ 0:10 1:20 1:10 1:5 1:2 1:1
DMAVFL 0.1076 0.0315 0.0152 0.0186 0.0188 0.0172
sync 0.2311 0.0240 0.0137 0.0166 0.0178 0.0189
@@
From Table 5 we can see that significant data leakage can occur, even in the absence of auxiliary data with the same distribution of the training set (i.e., with out-of-distribution auxiliary data). This observation is highlighted by a reconstruction error of 0.1076$\sim$0.2311, when the adversary uses FashionMNIST data to reconstruct images from MNIST. We also observe that for both DMAVFL and sync, the reconstruction error decreases quickly as $|D_{aux}|$ becomes relatively larger. However, after the ratio of $|D_{aux}|:|D_{train}|$ reaches 1:10, increasing auxiliary dataset size results in minor improvements in reconstruction error. This suggests that there is a threshold beyond which additional auxiliary data does not substantially improve the model’s ability to reconstruct target features, pointing to a diminishing return on the utility of expanding the auxiliary set.
7.4. Effect of the number of target clients
@@
Our proposed methods, DMAVFL and sync, are designed to handle scenarios involving multiple target clients. To assess the impact of varying the number of target clients on data reconstruction performance, we conduct experiments on the IOT dataset with the number of target clients ranging from 1 to 5. Consistent with the setup in the main experiments, 50% of the features are equally distributed among the target clients.
Table 6. Reconstruction errors for different numbers of target clients.
@@
DMAVFL 1.2789 1.5801 1.5986 1.6718 1.9820
sync 1.2324 1.6447 1.7597 1.8660 2.2635
The results presented in Table 6 indicate that an increase in the number of target clients, each sharing a smaller portion of the features, leads to diminished reconstruction performance. Specifically, the reconstruction error increases from 1.2789 with a single target client to 1.9828 when there are five target clients. This degradation in performance can be attributed to the distribution of features among more clients, each employing a separate bottom model to learn their respective features. This setup tends to cause the loss of crucial feature correlation information during the learning process. Moreover, as the encoder is trained on the complete set of features and is hence able to capture more feature correlations, the increased disparity between the encoder and the target models’ understanding of the feature correlation further exacerbates the challenge of accurate data reconstruction.
8. Evaluations on other defenses
While DMAVFL is designed and demonstrated to penetrate detection-based defenses of SplitGuard and Gradient Scrutinizer, we also evaluate its effectiveness against general defensive methods for privacy protection in VFL and SL. Note that unlike SplitGuard and Gradient Scrutinizer, these methods essentially introduce perturbations to the communicated embeddings or gradients, severely degrading the performance of honest training. Specifically, we consider three such defenses: Nopeek (Vepakomma et al., 2020), obfuscation with random noise (He et al., 2020), and differential privacy (DP) (Dwork, 2006; Abadi et al., 2016; Wei et al., 2020).
@@ 8. Evaluations on other defenses
(6) $\mathcal{L}_{N}$ $=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}(\alpha\cdot DCOR(\boldsymbol{x}_{j,n},f_{n}(\boldsymbol{x}_{j,n}))$
@@
We challenge Nopeek by deploying DMAVFL and sync against it with varying $\alpha$ values from 0.1 to 1.0 on the MNIST dataset. The results in Table 7 show that when $\alpha<0.9$, Nopeek fails to thwart our attacks. As $\alpha$ approaches 0.9 and beyond, Nopeek begins to impede the attack effectiveness. Notably, even at $\alpha=0.95$, DMAVFL still achieves a small reconstruction error of 0.0280. Setting $\alpha=1.0$ effectively defends the attack. However, this completely replaces the original VFL task with minimizing the $DCOR$ loss, which results in an accuracy of merely 69.03%.
Table 7. Reconstruction error against Nopeek.
@@
Obfuscation with random noise. Perturbation methods, such as adding noise to transmitted embeddings, are commonly employed in ML to safeguard privacy. Following the approach outlined in (He et al., 2020), we introduce noise with standard deviations $\sigma$ ranging from 0.2 to 0.8 into the embeddings sent by target clients. This defense is tested against DMAVFL and sync on the MNIST dataset, and the results are reported in Table 8. The results suggest that higher noise levels more effectively impede the attacks. This defense is more effective for sync, as its DAC learns a dynamic distribution of the encoding embedding during training, and is more susceptible to fluctuations caused by noise. Nonetheless, our attacks manage to reconstruct informative features even under significant noise, for instance, DMAVFL achieving a reconstruction error of 0.0472 at $\sigma=0.7$. It is also important to note that adding larger noise adversely affects the honest training accuracy.
Table 8. Reconstruction error against obfuscation.
@@
Differential privacy (DP). DP provides a mathematical framework for quantifying privacy guarantees for ML models. In our attacks, as the gradients returned from the active client may be maliciously generated, we propose to apply $\epsilon$-DP mechanism, which involves adding Laplacian noise to received gradients, to disrupt the malicious training. We test varying levels of privacy by adjusting the parameter $\epsilon$ from 0.1 to 10, where a smaller $\epsilon$ implies stronger privacy due to higher noise levels. The average reconstruction errors of DMAVFL and sync on MNIST are shown in Table 9, indicating that DP is most effective at an $\epsilon$ of 0.1, successfully impeding our proposed attacks. However, as $\epsilon$ increases, its defensive effectiveness decreases. At $\epsilon=10$, the noise assists the attack, potentially by making the model updates more robust, thereby facilitating the embedding distribution transfer. It’s important to note that while DP helps to protect privacy, it also hurts the accuracy of honest training.
Table 9. Reconstruction error against DP. $\epsilon=+\infty$ corresponds to the scenario without DP.
@@
9. Conclusion
We introduce DMAVFL, a novel defense-penetrating malicious data reconstruction attack for vertical federated learning (VFL). DMAVFL innovatively incorporates training of a discriminator with auxiliary classifier (DAC) to integrate label information into the attack process, which not only makes the malicious training indistinguishable from honest training, but also substantially improves data reconstruction performance. Through visualization and quantification, we demonstrated the effectiveness of DAC compared to traditional discriminators, showcasing its superior capability in embedding transfer. Through extensive experiments, we empirically demonstrated that DMAVFL successfully evades various defensive methods and outperforms all existing data reconstruction attacks. While traditional defense methods based on embedding/model perturbations are shown to effectively defend DMAVFL, it is at the cost of serious degradation on the honest VFL task’s performance. This calls for future research to develop effective defenses against malicious data reconstruction, with minimal impact on the honest task.
Appendix A DMAVFL_sync algorithm
We detail the steps of DMAVFL_sync in Algorithm 3.

### v2

@@ VI-B Baselines and detections
We also implement SOTA detection methods, SplitGuard (SG) and Gradient Scrutinizer (GS) on malicious attacks, to assess reconstruction performance under these defenses.
@@
Gradient Scrutinizer (GS): Evaluates the detection score of received gradients in each round to identify potential attacks. Due to instability in detection scores leading to frequent misjudgments, we adopt an average score strategy to improve reliability and set the threshold to 0.8 in our experiments. Similarly, when the average detection score falls below the detection threshold, an attack is detected and the bottom model update stop.
@@
In our experimental results (Table I, II, and III), we append the GS or SG to the names of baseline methods to indicate their evaluation under detection mechanisms, SG or GS.
VI-C Data reconstruction results
@@
Overview: As shown in Table I, II, and III, for all datasets, our proposed URVFL and URVFL_sync consistently outperform all other approaches in achieving the lowest reconstruction errors and embedding cosine distance. Our methods also achieve the best reconstruction performance measured by PSNR and SSIM compared with other baselines in image datasets. This underscores the effectiveness of our techniques and the advantage of embedding distribution transfer using DAC. For the Tiny imagenet, URVFL_sync achieves a reconstruction error of 0.0699, PSNR 17.5381, SSIM 0.3053, and cosine distance of 0.2139, significantly outperforming FSHA, which has a reconstruction error of 0.1025, PSNR 15.9321, SSIM 0.2205, and cosine distance of 0.3056.
@@
Our methods demonstrate robustness against detection mechanisms, maintaining consistent performance and circumventing both detection methods malicious attacks, i.e., URVFL and URVFL_sync is not detected by SG and GS. In sharp contrast, the performance of AGN and FSHA become worse when faced with the detections, and even completely failed under some circumstances. Specifically, AGN’s reconstruction error increases dramatically from 0.3786 to 1.9577, and FSHA’s from 0.1025 to 0.4797 for the Tiny imagenet dataset. This demonstrates the stealthiness and adaptability of URVFL and URVFL_sync compared to other methods.
An observation from the results is that a smaller embedding MSE distance does not guarantee a more accurate reconstruction or a more similar embedding distribution, as evidenced by GIA in the Credit dataset and PCAT in the RT_IOT2022 dataset, which has the smallest embedding MSE distance but significantly higher cosine embedding distance and reconstruction error compared to our methods. While minimizing the embedding MSE distance helps reconstruction, it does not guarantee a more similar embedding distribution, as metrics like cosine distance better capture the similarity between distributions. Generally, the decoder is trained to reconstruct data from the embedding distribution. Therefore, even if our methods have a higher embedding MSE distance, as long as the target model generates a similar embedding distribution, the decoder can still work effectively.
@@
Additionally, a smaller embedding cosine distance does not necessarily correspond to a lower reconstruction error in URVFL and URVFL_sync. For example, for the MNIST and RT_IOT2022 dataset, while URVFL_sync achieves a lower reconstruction error than URVFL, it exhibits a slightly higher embedding cosine distance. This can be attributed to the synchronous training of all model components in URVFL_sync, which promotes dynamic learning and results in model diversity. Despite the higher average embedding distance, the decoder in URVFL_sync is more effective at reconstructing features, effectively compensating for the increased cosine distance.
@@
We do not record AGN-GS and FSHA-GS in the Credit dataset. It is because GS doesn’t work in the Credit dataset, which can also be observed in Figure 6(j), caused by the limited data features (23 features) and imbalanced label distribution (77% label 0 and 23% label 1).
@@
Visualization of the reconstruction: Table IV visualizes the reconstruction performance of the proposed URVFL, alongside the malicious attack baselines, AGN and FSHA, for the CIFAR-10 dataset. We also report the visualization result of Tiny imagenet in Table XI in Appendix C. The results clearly demonstrate that URVFL produces more accurate and distinct reconstructions compared to the baselines. While AGN and FSHA yield coarser results, URVFL maintains high-quality reconstruction even under SG detection. In contrast, both FSHA and AGN struggle to reconstruct informative pixels effectively when faced with detection mechanism SG.
@@
TABLE IV: Visualization of reconstructed features on the CIFAR-10 dataset with and without the detection SG. Except for the original image, the right side of each image is the reconstructed image. The last column is the average reconstruction error.
@@
SG and GS detection results: Figure 6 illustrates the detection scores for URVFL and URVFL_sync, which consistently exceed the detection thresholds, aligning indistinguishably with honest training scenarios. Although in the CIFAR-10 and Tiny imagenet dataset, URVFL’s detection score is lower than the threshold at the beginning, its average score remains above the threshold, ensuring stealthiness against target clients. In contrast, malicious attack baselines, AGN and FSHA, consistently maintain scores below the detection thresholds of SG and GS in the MNIST and CIFAR-10 datasets, rendering them detectable.
@@
Embedding distances across attack iterations: Figure 7 visualizes the changes in embedding distance during the attack for our methods compared to FSHA on the CIFAR-10 dataset. The figure clearly demonstrates that our methods steadily reduce the embedding distance throughout the attack, resulting in lower overall embedding distances and effectively aligning the embedding distribution with the target model. In contrast, FSHA exhibits less stability, despite also showing a decreasing trend in embedding distance.
Furthermore, we observe that the SG score for a single iteration of our methods falls below the threshold within the first 10 iterations in Fig. 6, whereas FSHA remains below the threshold throughout the entire training process, leading to the detection of the attack. This difference may be attributed to the more drastic changes in embedding distance during our training process. As illustrated in Figure 7, our methods cause rapid changes in embedding distance initially, which then slow down, while FSHA maintains a more consistent fluctuations in embedding distance throughout the attack.
@@
(c) SG scores in Tiny.
(d) SG scores in IOT.
(e) SG scores in Credit.
(f) GS scores in MNIST.
(g) GS scores in CIFAR-10.
(h) GS scores in Tiny.
(i) GS scores in IOT.
(j) GS scores in Credit.
Figure 6: Detection scores of SG and GS on malicious attacks and the honest training across four datasets. We abbreviate Tiny imagenet dataset to Tiny and RT_IOT2022 dataset to IOT.
Figure 7: Embedding MSE and cosine distances across attack iterations using FSHA, URVFL, and URVFL_sync.
VI-D Ablation Study
This section explores the robustness of URVFL and URVFL_sync (abbreviated as sync) by evaluating their performance under various parameter settings. These settings include changes in the adversary’s feature size, use of heterogeneous encoder models, variations in auxiliary dataset size, and differences in the number of target clients. Except for the variables specifically altered for this study, the model setups and implementation environment remain consistent with those described in the main experiments.
VI-D1 Effect of adversary’s feature size
To explore how the size of the feature set held by the adversary affects the reconstruction performance of URVFL and sync, we conduct experiments targeting the reconstruction of 30% of the features while varying the proportion of features held by the adversary from 0% to 70%. At 0% feature possession, the scenario reduces to split learning where no features or bottom model are held by the adversary. These experiments are performed on the Credit and RT_IOT2022 (abbreviated as IOT) datasets, with the results summarized in Figure 9.
Figure 8: Reconstruction error under various adversary’s feature portions.
Figure 9: Reconstruction error for different $|D_{aux}|$ relative to $|D_{train}|.$
Consistent with our expectations, increasing the proportion of features held by the adversary leads to less reconstruction error. As illustrated in Figure 9, the reconstruction error in the Credit dataset using URVFL decreases from 0.8315 to 0.5189 as the feature possession increases from 0% to 70%. This trend is observed for both datasets and methods.
VI-D2 Effect of encoder’s complexity
@@
TABLE V: Reconstruction error and average embeddings distance using different encoder’s model.
Encoder Metric URVFL sync
@@
Table V clearly illustrates that the adversary achieves the lowest reconstruction error and the smallest embedding cosine distance when using an encoder the same as the target model, as observed for both URVFL and sync. This optimal alignment underscores the importance of model congruence in enhancing the effectiveness of embedding distribution transfer and data reconstruction. Conversely, when the encoder structure deviates from that of the target model, specifically when switching from a CNN to an MLP, the reconstruction performance significantly deteriorates, with reconstruction error increasing by an order of magnitude (from 0.0302 to 0.2545). This dramatic drop highlights the challenges of using less compatible models for data reconstruction attack. Moreover, the results indicate that a complex encoder, featuring a higher number of residual blocks, outperforms a simpler encoder configuration. The complex encoder not only enhances the decoder’s ability to reconstruct the data accurately but also learns a more informative embedding. This suggests that increasing the encoder’s complexity, when the target model is unknown, can lead to better performance in both embedding distribution transfer and data reconstruction.
VI-D3 Effect of auxiliary dataset
@@
In the main experiments, we set the size of the adversary’s auxiliary dataset, $|D_{aux}|$, relative to the training set size, $|D_{train}|$, at a ratio of $1:10$. To explore the impact of varying auxiliary dataset sizes, we adjust the ratio of $|D_{aux}|:|D_{train}|$ from $0:10$ to $1:1$ in the MNIST and CIFAR10 dataset. Specifically, at a ratio of 0:10, the auxiliary dataset of MNIST and CIFAR-10 comprise the FashionMNIST [42] and the downsampled Tiny imagenet, respectively, introducing a distinct distribution from the training set. This setup aims to test the adaptability of the reconstruction process to changes in the auxiliary dataset’s distribution.
@@
From Figure 9 we can see that significant data leakage can occur, even in the absence of auxiliary data with the same distribution of the training set (i.e., with out-of-distribution auxiliary data). This observation is highlighted by a reconstruction error of 0.1076$\sim$0.2311, when the adversary uses FashionMNIST data to reconstruct images from MNIST, and 0.1696$\sim$0.1906 when the adversary uses Tiny imagenet data to reconstruct images from CIFAR-10. We also observe that for both URVFL and sync, the reconstruction error decreases quickly as $|D_{aux}|$ becomes relatively larger. However, after the ratio of $|D_{aux}|:|D_{train}|$ reaches 1:10, increasing auxiliary dataset size results in minor improvements in reconstruction error. This suggests that there is a threshold beyond which additional auxiliary data does not substantially improve the model’s ability to reconstruct target features, pointing to a diminishing return on the utility of expanding the auxiliary set.
VI-D4 Effect of the number of target clients
@@
Our proposed methods, URVFL and sync, are designed to handle scenarios involving multiple target clients. To assess the impact of varying the number of target clients on data reconstruction performance, we conduct experiments on the IOT dataset with the number of target clients ranging from 1 to 5. Consistent with the setup in the main experiments, 50% of the features are equally distributed among the target clients.
TABLE VI: Reconstruction errors for different numbers of target clients.
@@
URVFL 1.3821 1.5801 1.5986 1.6718 1.9820
sync 1.3277 1.6447 1.7597 1.8660 2.2635
The results presented in Table VI indicate that an increase in the number of target clients, each sharing a smaller portion of the features, leads to diminished reconstruction performance. Specifically, the reconstruction error increases from 1.3821 with a single target client to 1.9820 when there are five target clients. This degradation in performance can be attributed to the distribution of features among more clients, each employing a separate bottom model to learn their respective features. This setup tends to cause the loss of crucial feature correlation information during the learning process. Moreover, as the encoder is trained on the complete set of features and is hence able to capture more feature correlations, the increased disparity between the encoder and the target models’ understanding of the feature correlation further exacerbates the challenge of accurate data reconstruction.
VII Evaluations on defenses
While URVFL is designed and demonstrated to penetrate detection-based methods of SplitGuard and Gradient Scrutinizer, we also evaluate its effectiveness against general defensive methods for privacy protection in VFL and SL. Note that unlike SplitGuard and Gradient Scrutinizer, these methods essentially introduce perturbations to the communicated embeddings or gradients, severely degrading the performance of honest training. Specifically, we consider three such defenses: Nopeek [37], obfuscation with random noise [15], and differential privacy (DP) [5, 1, 40].
@@ VII Evaluations on defenses
$\mathcal{L}_{N}$ $=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}(\alpha\cdot DCOR(\boldsymbol{x}_{j,n},f_{n}(\boldsymbol{x}_{j,n}))$ (6)
@@
We challenge Nopeek by deploying URVFL and sync against it with varying $\alpha$ values from 0.1 to 1.0 on the MNIST dataset. The results in Table VII show that when $\alpha<0.9$, Nopeek fails to thwart our attacks. As $\alpha$ approaches 0.9 and beyond, Nopeek begins to impede the attack effectiveness. Notably, even at $\alpha=0.95$, URVFL still achieves a small reconstruction error of 0.0280. Setting $\alpha=1.0$ effectively defends the attack. However, this completely replaces the original VFL task with minimizing the $DCOR$ loss, which results in an accuracy of merely 69.03%.
TABLE VII: Reconstruction error against Nopeek.
@@
Obfuscation with random noise. Perturbation methods, such as adding noise to transmitted embeddings, are commonly employed in ML to safeguard privacy. Following the approach outlined in [15], we introduce noise with standard deviations $\sigma$ ranging from 0.2 to 0.8 into the embeddings sent by target clients. This defense is tested against URVFL and sync on the MNIST dataset, and the results are reported in Table VIII. The results suggest that higher noise levels more effectively impede the attacks. This defense is more effective for sync, as its DAC learns a dynamic distribution of the encoding embedding during training, and is more susceptible to fluctuations caused by noise. Nonetheless, our attacks manage to reconstruct informative features even under significant noise, for instance, URVFL achieving a reconstruction error of 0.0472 at $\sigma=0.7$. It is also important to note that adding larger noise adversely affects the honest training accuracy.
TABLE VIII: Reconstruction error against obfuscation.
@@
Differential privacy (DP). DP provides a mathematical framework for quantifying privacy guarantees for ML models. In our attacks, as the gradients returned from the active client may be maliciously generated, we propose to apply $\epsilon$-DP mechanism, which involves adding Laplacian noise to received gradients, to disrupt the malicious training. We test varying levels of privacy by adjusting the parameter $\epsilon$ from 0.1 to 10, where a smaller $\epsilon$ implies stronger privacy due to higher noise levels. The average reconstruction errors of URVFL and sync on MNIST are shown in Table IX, indicating that DP is most effective at an $\epsilon$ of 0.1, successfully impeding our proposed attacks. However, as $\epsilon$ increases, its defensive effectiveness decreases. At $\epsilon=10$, the noise assists the attack, potentially by making the model updates more robust, thereby facilitating the embedding distribution transfer. It’s important to note that while DP helps to protect privacy, it also hurts the accuracy of honest training.
TABLE IX: Reconstruction error against DP. $\epsilon=+\infty$ corresponds to the scenario without DP.
@@
VIII Potential detection of our attacks
In this section, we explore potential detection mechanisms tailored to identify our proposed attacks. Our attack integrates label information into the gradients, rendering simple observations of gradient differences across various labels ineffective for detection. Specifically, our approach not only classifies the corresponding labels of embeddings but also distinguishes between true and fake embeddings generated by either the local encoder or the target model. Consequently, we hypothesize that the increased number of labels and the incorporation of a GAN-like structure will influence the distribution of gradient norms. To investigate this, we analyze the distribution of gradients’ L2 norms across both tabular and image datasets.
(a) Embedding gradients norm in Credit dataset.
(b) Embedding gradients norm in IOT dataset.
(c) Embedding gradients norm in CIFAR-10 dataset.
(d) Embedding gradients norm in Tiny imagenet dataset.
Figure 10: Distribution of Embedding Gradients’ L2 Norms in Honest Training versus Our URVFL Attack.
Fig.10 illustrates a clear distinction in the distribution of embedding gradients’ L2 norms between honest training and our URVFL attack. In tabular datasets, the gradient norms under attack are sparsely distributed with a higher average value, whereas, in honest training, they concentrate around lower values. Conversely, for image datasets, where the pattern reverses, honest training results in sparsely distributed gradients with higher average norms, while the URVFL attack leads to concentrated gradients with lower average norms. This discrepancy is likely attributable to inherent differences in data characteristics and model architectures between image and tabular datasets.
Despite the varying gradient norm distributions between tabular and image datasets, a consistent and significant difference remains evident between honest training and the URVFL attack across both data types. Leveraging this insight, a passive client can enhance detection capabilities by adopting a two-step strategy. Initially, the passive client collects publicly available comprehensive training data and conducts local training for several hundred iterations to establish a baseline distribution of embedding gradients’ L2 norms under honest training conditions. This baseline serves as a reference point for normal gradient behavior.
During subsequent training sessions, the passive client continuously records the distribution of embedding gradients. By comparing the observed gradient distribution against the pre-established honest training distribution, the client can identify significant deviations indicative of an attack. If the gradient norms exhibit patterns that deviate substantially from the baseline, such as increased sparsity and higher average norms in tabular datasets or the opposite pattern in image datasets, the passive client can infer the presence of an attack. Upon detecting such discrepancies, the passive client can stop updating the local bottom model, thereby preventing further compromise and safeguarding the integrity and privacy of the training process.
IX Conclusion
We introduce URVFL, a novel undetectable malicious data reconstruction attack for vertical federated learning (VFL). URVFL innovatively incorporates training of a discriminator with auxiliary classifier (DAC) to integrate label information for malicious gradient generation, which not only makes the malicious training indistinguishable from honest training, but also substantially improves data reconstruction performance. Through visualization and quantification, we demonstrated the effectiveness of DAC compared to traditional discriminators, showcasing its superior capability in embedding transfer. Through extensive experiments, we empirically demonstrated that URVFL successfully evades SOTA detection methods and outperforms all existing data reconstruction attacks. While traditional defense methods based on embedding/model perturbations are shown to effectively defend URVFL, it is at the cost of serious degradation on the honest VFL task’s performance. This calls for future research to develop effective defenses against malicious data reconstruction, with minimal impact on the honest task.
Acknowledgment
The work of Songze Li is in part supported by the National Nature Science Foundation of China (NSFC) Grant 62106057, and the Fundamental Research Funds for the Central Universities (Grant No. 2242024k30059).
Appendix A URVFL_sync algorithm
We detail the steps of URVFL_sync in Algorithm 3.